3-1 ka harp fun 9.3 ka and sainp all have the Computed value: 18242 Submitted: Ihandry, 4 is moving N Stars contant ka hang from
Answer #1

1)

x = stretch of the spring = 0.12 m

k = spring constant

m = mass hanged = 3.1 kg

using equilibrium of force in vertical direction

spring force = weight of mass hanged

k x = mg

k (0.12) = 3.1 x 9.8

k = 253.2 N/m

3)

for the lower spring

spring force = (m2 + m3) g

k x = (9.3 + 6.2) (9.8)

253.2 x = (9.3 + 6.2) (9.8)

x = 0.6 m